The lenght and width of a rectangle are in the ratio 3:2. If the length is increased by 20 and teh width is decreased by 5, the ratio of the length to the width becomes 2:1. Find teh length and width of this rectangle. : The lenght and width of a rectangle are in the ratio 3:2. If the length is increased by 20 and teh width is decreased by 5, the ratio of the length to the width becomes 2:1. Find teh length and width of this rectangle.
Answer by mangopeeler07(423)

L/w=3/2

Cross multiply
3w=2L
Divide by 3
w=2L/3

L+20/w-5=2/1
Cross multiply
2w-10=L+20

Plug in w
2(2L/3)-10=L+20

Distribute
4L/3-10=L+20

Subtract 4/3L from both sides
-10=-1/3L+20

Subtract 20 from both sides
-30=-1/3L

Multiply both sides by -3
90=L

90/w=3/2
Cross multiply
180=3w
60=w

Answer:
Length=90
Width=60